# Reinforced Latent Reasoning for LLM-based Recommendation

### v1

@@ 4 Experiments
In this section, we conduct experiments to answer the following research questions: RQ1: How does the overall recommendation performance of our proposed LatentR3 model compare to existing recommendation methods? RQ2: Where do the improvements of LatentR3 come from? RQ3: How do different design choices and reasoning hyperparameters affect the effectiveness of our method?
@@ 4.1 Experimental Settings
In line with existing works, we conduct experiments on multiple domain-specific datasets from the Amazon review corpus, specifically focusing on three sub-datasets: Toys, CDs, and Games. Following the standard preprocessing procedure, we apply 5-core filtering to ensure data quality. Additionally, we adopt several temporal filtering strategies to maintain a manageable dataset size for RL training while preserving data quality—such as ensuring each dataset contains more than 5,000 items. More detailed preprocessing steps and dataset statistics are provided in Appendix A.
@@
We consider the top-N recommendation task. To evaluate the effectiveness of our method, we compare it against both 1) representative traditional sequential recommendation models, including Caser (Tang and Wang, 2018), GRU4Rec (Hidasi and Karatzoglou, 2018), and SASRec (Kang and McAuley, 2018); and 2) LLM-based recommendation approaches, including AlphaRec (Sheng et al., 2025), BIGRec (Bao et al., 2025), and D3 Bao et al. (2024). AlphaRec is a state-of-the-art (SOTA) method that leverages LLM-generated embeddings to enhance the recommendation model, while D3 is an SOTA method that directly leverages LLMs as generative recommenders. More details about the baseline can be found in Append B. For our LatentR3, we implement it on both BIGRec and D3. Notably, existing LLM reasoning methods for recommendation cannot be selected for comparison because: (1) they are not designed for top-N recommendation tasks, and (2) their training relies on explicit CoT supervision, which is difficult to obtain in our setting. Similarly, we exclude general-domain latent reasoning methods for comparison, as they still depend on CoT data for learning. However, in our ablation study, we will include a variant without RL training, which can be viewed as an adaptation of the general latent reasoning method to our setting. The implementation details, including hyperparameter tuning, can be found in Appendix C.
@@
We evaluate the top-N recommendation effectiveness using Hit Ratio (HR@N) and Normalized Discounted Cumulative Gain (NDCG@N), with N set to 5 and 10. For brevity, we denote HR@5 and NDCG@5 as H@5 and N@5, respectively, in tables and figures; the same applies to the ‘@10’ cases.
@@ 4.2 Main Results (RQ1)
Table 1: Top-N recommendation performance comparison between our LatentR3 and baseline methods. “RI” denotes the relative improvement of our best-performing version (latentR3 applied to D3) over each baseline across all datasets and metrics. By default, the comparison is made against each column. For the BIGRec column specifically, the RI is calculated using LatentR3 applied to BIGRec. The best results are bold.
@@
Caser GRU4Rec SASRec AlphaRec BIGRec +LatentR3 D3 +LatentR3
Toys H@5 0.0251 0.0417 0.0601 0.0579 0.0701 0.0821 0.0830 0.0898
H@10 0.0384 0.0564 0.0760 0.0893 0.0931 0.1107 0.1026 0.1152
N@5 0.0170 0.0305 0.0458 0.0347 0.0508 0.0600 0.0610 0.0670
N@10 0.0214 0.0352 0.0510 0.0448 0.0582 0.0693 0.0674 0.0752
CDs H@5 0.0469 0.0481 0.0841 0.0479 0.0757 0.0934 0.1122 0.1137
H@10 0.0689 0.0669 0.1054 0.0774 0.0929 0.1160 0.1272 0.1327
N@5 0.0312 0.0365 0.0622 0.0278 0.0616 0.0754 0.0906 0.0915
N@10 0.0382 0.0425 0.0691 0.0373 0.0672 0.0826 0.0955 0.0977
Games H@5 0.0324 0.0322 0.0416 0.0558 0.0461 0.0580 0.0608 0.0716
H@10 0.0538 0.0517 0.0633 0.0893 0.0709 0.0870 0.0860 0.1006
N@5 0.0211 0.0207 0.0280 0.0397 0.0334 0.0413 0.0423 0.0507
N@10 0.0280 0.0270 0.0350 0.0515 0.0414 0.0506 0.0505 0.0601
RI 170.8% 121.0% 52.3% 77.5% 21.8% - 10.4% -
Table 1 presents the overall performance comparison between our method and the baselines, including both traditional (Caser, GRU4Rec, SASRec) and LLM-based (AlphaRec, BIGRec, D3) approaches. From the results, we draw the following conclusions:
@@
  • Focusing on the comparison among LLM-based methods, those that use LLMs as recommenders (BIGRec and D3) outperform the method that leverages LLMs to enhance traditional models (AlphaRec) on the Toys and CDs datasets, but underperform on the Games dataset. After applying our LatentR3, the performance of both BIGRec and D3 improves significantly, with BIGRec achieving a relative improvement (RI) of 21.8% and D3 achieving an RI of 10.4%. Furthermore, in most cases, these enhanced models surpass AlphaRec. These results demonstrate that 1) incorporating latent reasoning substantially improves the performance of the LLM-based methods that leverage LLMs as the recommender; 2) our latent reasoning method can be applied to different existing LLM-based methods.
  • The best-performing LLM-based baselines generally outperform the strongest traditional baselines, consistent with findings in prior work. This advantage is largely attributed to the superior capabilities of LLMs, such as their ability to model rich world knowledge.
@@ 4.3.1 Performance over Popular and Unpopular Items
For latent reasoning, it is difficult to directly demonstrate why it leads to performance improvements, as can be done with explicit CoT. Therefore, we consider providing indirect evidence. Intuitively, reasoning should be more beneficial for the more challenging aspects of recommendation, where it is expected to deliver greater gains. Long-tail (i.e., unpopular) items are typically more difficult to recommend accurately, while existing methods already perform well on popular items. To test this hypothesis, we examine whether our method yields greater improvements on unpopular items. Specifically, we compare the relative performance gains of our method over BIGRec across items with different popularity levels. The results are shown in Figure 2 (see Appendix D for more detailed results and experimental settings). As shown, our method achieves significantly larger improvements on unpopular items, suggesting that the incorporation of reasoning is particularly beneficial in more challenging recommendation scenarios.
@@ 4.3.2 Ablation Study
Our method features key innovations in both the model architecture and the learning algorithm. To assess the contribution of each component, we conduct ablation studies on both aspects. Specifically, we evaluate the following variants of our method: 1) w/o Reasoning, which removes the latent reasoning component entirely. 2) w/o LatentRATT, which omits the LatentRATT layer and directly uses the final hidden states from the LLM to construct the latent reasoning token; 3) w/o RL, which trains the reasoning model using only supervised fine-tuning (SFT), without RL; 4) w/o Batch Advantage, which replaces our batch-based advantage estimation with the original GRPO setting that computes advantages with a group average reward. We compare our default method with these variants on the Toys and CDs datasets. Notably, regarding the reward design, we would study it later when compared to the original GRPO.
@@ 4.3.3 Influence of Reasoning Length
[…]
@@ 4.3.4 Comparison with the Original GRPO
We also compare our modified GRPO with the original GRPO, which relies on generating complete answers for reward computation for group sampling, to evaluate its effectiveness. To improve training efficiency, we propose using the model’s perplexity (PPL) on the ground-truth answer as a proxy for the reward. This yields a continuous reward signal and motivates adjustments to the advantage computation. To assess the impact of these modifications, we implement a variant of our method using the original GRPO to implement the RL. The comparison results on the CDs dataset are shown in Figure 4, with additional results and implementation details provided in Appendix F. In terms of performance, our method achieves results comparable to the variant using original GRPO—slightly outperforming it on the CDs dataset while performing slightly worse on the Toys dataset (see Appendix F). In terms of training efficiency, our method is significantly faster, reducing the training cost to approximately 1/5 of that of the original GRPO-based method. These findings validate the rationale and effectiveness of our proposed modifications.
@@ 5 Related Work
Our work relates to LLM-based recommendation and latent reasoning. Due to space constraints, we briefly discuss LLM-based recommendation and its associated reasoning methods here, with a more comprehensive review available in Appendix G. The recent widespread success of LLMs has sparked growing interest in their application to recommendation, giving rise to three main LLM-based recommendation paradigms: (1) in-context learning (Gao et al., 2023; Sun et al., 2024; Wang et al., 2021), (2) agent-based frameworks (Wang et al., 2023; Zhang et al., 2024a, b), and (3) fine-tuning-based methods (Bao et al., 2025; Zhang et al., 2024c). Among these, the fine-tuning-based solution has received the most attention due to its ability to more effectively align LLMs with the task, and our work falls into this category. In this direction, only a few studies have explored reasoning, with existing works primarily aiming to enhance explicit chain-of-thought (CoT) reasoning via fine-tuning, such as RecSAVER (Tsai et al., 2024), ReasoningRec (Bismay et al., 2024), Reasoning4Rec (Fang et al., 2025), and Exp3rt (Kim et al., ). However, these approaches rely heavily on CoT data, which is costly to obtain and often difficult to ensure in quality. Moreover, the use of explicit CoT reasoning typically incurs higher inference latency. In contrast, our work introduces latent reasoning for LLM-based recommendation, which eliminates the need for CoT supervision and enables more efficient inference. As for latent reasoning, only two concurrent works (Tang et al., 2025; Zhang et al., 2025) have explored it, but neither focuses on LLM-based recommendation, and our approach to latent reasoning is fundamentally different and specifically tailored for LLM-based recommendation.
@@ 6 Conclusion
Due to current computational resource constraints, our experiments are limited to relatively small datasets. In the future, with access to greater computational resources, we plan to explore the effectiveness of our method on larger-scale datasets. Additionally, our method currently faces challenges when generating long latent reasoning sequences. Future work will focus on developing more effective and efficient approaches to learn latent reasoning with extended reasoning lengths. Last but not least, compared to explicit CoT, latent reasoning offers less interpretability. We intend to investigate ways to improve the explainability of latent reasoning, which could be crucial for its broader adoption and practical use.
@@ Appendix A Dataset Pre-processing
Our experiments were conducted on the Amazon review dataset, with primary evaluations performed on three domain-specific subsets: Toys, CDs, and Games. Following established methodology, we preprocessed each subset to achieve 5-core data integrity. To address the computational demands of LLM training while maintaining dataset validity, we implemented a dynamic temporal partitioning strategy: We iteratively process the dataset within a sliding time window, starting from October 2017 to October 2018. After applying 5-core filtering, if the resulting item count falls below 5,000, the start time is shifted back by 3 months (e.g., to July 2017). This adjustment repeats until the processed dataset contains more than 5,000 items, with the end time fixed at October 2018 throughout the process. Key dataset statistics are summarized in the accompanying table3. To align with real-world scenarios, we adopt a temporal split of the processed datasets into training, validation, and testing sets based on timestamps, following an 8:1:1 ratio. In alignment with comparative baseline requirements across all experimental conditions, we standardized the maximum sequence length at 10 for consistency in temporal pattern analysis.
@@
Appendix B Compared Methods
We consider the top-N recommendation task. To evaluate the effectiveness of our method, we compare it against both representative traditional sequential recommendation models and state-of-the-art (SOTA) LLM-based recommendation approaches:
    Caser (Tang and Wang, 2018): This is a well-known sequential recommendation approach that utilizes Convolutional Neural Networks (CNNs) to capture sequential patterns and model user preferences.
    GRU4Rec (Hidasi and Karatzoglou, 2018): This is another widely recognized method that leverages Gated Recurrent Units (GRUs) to encode sequential patterns and model user preferences.
    SASRec (Kang and McAuley, 2018): This is a highly representative sequential recommendation method that utilizes a self-attention network to model user preferences, with an architecture resembling that of the decoder-only Transformer model.

### v2

@@ 4 Experiments
In this section, we conduct experiments to answer the following research questions: RQ1: How does the overall recommendation performance of our proposed LatentR3 method compare to existing recommendation methods? RQ2: Where do the improvements of LatentR3 come from? RQ3: How do different design choices and reasoning hyperparameters affect the effectiveness of our method?
@@ 4.1 Experimental Settings
In line with prior LLM-based recommendation studies, we evaluate our method on several Amazon domain-specific datasets, including Toys, CDs, Games and Instruments. We adopt the standard preprocessing procedure that preserves chronological order for data filtering and splitting, with detailed steps and dataset statistics provided in Appendix A.
@@
We consider the top-$N$ recommendation task and compare our LatentR3 with: (1) traditional sequential models, including Caser (Tang & Wang, 2018), GRU4Rec (Hidasi & Karatzoglou, 2018), and SASRec (Kang & McAuley, 2018); and (2) LLM-based approaches, including Base (direct LLM recommendation), COT (Tsai et al., 2024), AlphaRec (Sheng et al., 2025), BIGRec (Bao et al., 2025), and D3 (Bao et al., 2024). AlphaRec is a state-of-the-art embedding-enhanced model, while D3 is a state-of-the-art generative recommender. We implement LatentR3 on both BIGRec and D3. Details on baselines and implementation are provided in Appendix B. Here, for explicit reasoning, we only include the direct CoT baseline. Other reasoning methods for recommendation mainly focus on rating prediction and require CoT supervision, so we exclude them. We have tried learning explicit reasoning without CoT supervision via RL, but it performed poorly (Appendix C), and we thus also omit it here. We also exclude general latent reasoning methods since they rely on CoT data, but include a no-RL variant in ablation study as their adaptation to our setting.
@@
We evaluate the top-N recommendation effectiveness using Hit Ratio (HR@N) and Normalized Discounted Cumulative Gain (NDCG@N), with N set to 5 and 10. For brevity, we denote HR@5 and NDCG@5 as H@5 and N@5, respectively, similarly for ‘@10’ cases.
@@ 4.2 Main Results (RQ1)
Table 1: Top-NN recommendation performance of LatentR3 versus baselines. "RI." indicates the relative improvement of LatentR3 on D3 over each baseline (for the BIGRec column, "RI." uses LatentR3 on BIGRec) across all datasets and metrics. Bold values denote the best results. The analysis of statistical significance is reported in the Appendix D, and our results are statistically significant.
@@
Metrics Caser GRU4Rec SASRec Base COT AlphaRec BIGRec +LatentR3 D3 +LatentR3
Toys H@5 0.0251 0.0417 0.0601 0.0203 0.0261 0.0579 0.0701 0.0821 0.0830 0.0898
H@10 0.0384 0.0564 0.0760 0.0359 0.0496 0.0893 0.0931 0.1107 0.1026 0.1152
N@5 0.0170 0.0305 0.0458 0.0128 0.0153 0.0347 0.0508 0.0600 0.0610 0.0670
N@10 0.0214 0.0352 0.0510 0.0178 0.0229 0.0448 0.0582 0.0693 0.0674 0.0752
CDs H@5 0.0469 0.0481 0.0841 0.0195 0.0302 0.0479 0.0757 0.0934 0.1122 0.1137
H@10 0.0689 0.0669 0.1054 0.0252 0.0406 0.0774 0.0929 0.1160 0.1272 0.1327
N@5 0.0312 0.0365 0.0622 0.0148 0.0213 0.0278 0.0616 0.0754 0.0906 0.0915
N@10 0.0382 0.0425 0.0691 0.0167 0.0246 0.0373 0.0672 0.0826 0.0955 0.0977
Games H@5 0.0324 0.0322 0.0416 0.0236 0.0120 0.0558 0.0461 0.0580 0.0608 0.0716
H@10 0.0538 0.0517 0.0633 0.0311 0.0194 0.0893 0.0709 0.0870 0.0860 0.1006
N@5 0.0211 0.0207 0.0280 0.0190 0.0082 0.0397 0.0334 0.0413 0.0423 0.0507
N@10 0.0280 0.0270 0.0350 0.0214 0.0105 0.0515 0.0414 0.0506 0.0505 0.0601
Instruments H@5 0.0564 0.0630 0.0708 0.0296 0.0261 0.0564 0.0807 0.0882 0.0848 0.0920
H@10 0.0627 0.0692 0.0758 0.0411 0.0452 0.0640 0.0879 0.0941 0.0907 0.0973
N@5 0.0781 0.0766 0.0793 0.0154 0.0135 0.0813 0.0938 0.1029 0.0984 0.1066
N@10 0.0977 0.0960 0.0950 0.0192 0.0199 0.1051 0.1158 0.1214 0.1167 0.1229
RI 85.8% 67.8% 27.9% 266.8% 245.9% 38.8% 17.0% - 8.4% -
Table 1 presents the overall performance comparison between our method and the baselines, including both traditional (Caser, GRU4Rec, SASRec) and LLM-based (w/o tuning: Base, COT; w/ tuning: AlphaRec, BIGRec, D3) approaches. From the results, we draw the following findings:
@@
  • Focusing on the comparison among the LLM-based methods with recommendation-specific tuning, those that use LLMs as recommenders (BIGRec and D3) outperform the method that leverages LLMs to enhance traditional models (AlphaRec) on the Toys, CDs, and Instruments, but underperform on Games. After applying our LatentR3, the performance of both BIGRec and D3 improves significantly, with BIGRec achieving a relative improvement (RI) of 17.0% and D3 achieving an RI of 8.4%. Furthermore, the enhanced models surpass AlphaRec. These results show that 1) incorporating latent reasoning substantially improves the performance of the LLM-based methods; 2) our latent reasoning method can be applied to different existing LLM-based methods.
  • Directly using LLMs for recommendation without tuning (Base) performs substantially worse than traditional methods. Adding explicit reasoning via CoT provides slight improvements but still largely falls short of traditional baselines. After tuning, the best-performing LLM-based baselines outperform traditional models in nearly all cases. These results underscore the importance of explicitly aligning LLMs—including their reasoning capabilities—with the recommendation task.
To further validate our method, Appendix E compares LatentR3 with BIGRec on a larger LLM, showing consistent gains.
Efficiency Comparison: Beyond accuracy, a big advantage of our method is that it needs only a few latent tokens (one in our experiments) to represent reasoning, greatly reducing inference latency. An empirical study on four datasets (Appendix F) shows that our method maintains a cost close to non-reasoning LLM baselines, whereas explicit COT incurs a substantial increase (e.g., a 25-fold increase on Toys and nearly a 30-fold increase on Games).
@@ 4.3.1 Performance over Popular and Unpopular Items
For latent reasoning, it is difficult to directly demonstrate why it leads to performance improvements, as can be done with explicit CoT. Therefore, we consider providing indirect evidence. Intuitively, reasoning should be more beneficial for the more challenging aspects of recommendation, where it is expected to deliver greater gains. Long-tail (i.e., unpopular) items are typically more difficult to recommend accurately, while existing methods already perform well on popular items. To test this hypothesis, we examine whether our method yields greater improvements on unpopular items. Specifically, we compare the relative performance gains of our method over BIGRec across items with different popularity levels. The results are shown in Figure 2 (see Appendix G for more results and experimental settings). As shown, our method achieves significantly larger improvements on unpopular items, suggesting that the incorporation of reasoning is particularly beneficial in more challenging recommendation scenarios.
@@ 4.3.2 Ablation Study
Our method features key innovations in both the model architecture and the learning algorithm. To assess the contribution of each component, we conduct ablation studies on both aspects. Specifically, we evaluate the following variants of our method: 1) w/o Reasoning, which removes the latent reasoning component entirely. 2) w/o LatentRATT, which omits the LatentRATT layer and directly uses the final hidden states from the LLM to construct the latent reasoning token; 3) w/o RL, which trains the reasoning model using only SFT without RL; 4) w/o Batch Advantage, which replaces our batch-based advantage estimation with the original GRPO setting that computes advantages with a group average reward. Notably, regarding the reward design, we would study it later when compared to the original GRPO. We conduct experiments on the Toys and CDs datasets.
@@ 4.3.3 Influence of Reasoning Length
We next investigate how the length of latent reasoning affects model performance by varying the number of generated latent tokens ($K$). Considering the cost of RL tuning, we just tune $K$ in {0,1,2}, and conducted experiments on Toys and CDs. The results in terms of NDCG@5 (N@5) and HR@5 (H@5) are shown in Figure 3. Results for other metrics and the detailed implementations can be found in Appendix H. As shown, increasing the reasoning length improves performance, demonstrating the potential to achieve even better results by increasing the reasoning length. However, the gain from $K\text{=}1$ to $K\text{=}2$ is much smaller than from $K\text{=}0$ to $K\text{=}1$, suggesting that a few latent tokens may be sufficient to capture most of the reasoning information.
@@ 4.3.4 Comparison with the Original GRPO
We also compare our LR-GRPO with the original GRPO, which relies on generating complete answers for reward computation for group sampling, to evaluate its effectiveness. To improve training efficiency, we propose using the model’s perplexity (PPL) on the ground-truth answer as a proxy for the reward. This yields a continuous reward signal and motivates adjustments to the advantage computation. To assess the impact of these modifications, we implement a variant of our method using the original GRPO. The comparison results on the CDs dataset are shown in Figure 4, with additional results and implementation details provided in Appendix I. In terms of performance, our method achieves results comparable to the variant using original GRPO—slightly outperforming it on CDs while performing slightly worse on Toys (see Appendix I). In terms of training efficiency, our method is significantly faster, reducing the training cost to approximately 1/4 of that of the original GRPO. These findings validate the rationale of our modifications.
@@ 5 Related Work
Our work relates to LLM-based recommendation and latent reasoning. Due to space limits, we briefly cover LLM-based recommendation and reasoning methods here; a full review is in Appendix J. Existing works include three main LLM-based recommendation paradigms: (1) in-context learning (Gao et al., 2023; Sun et al., 2024; Wang et al., 2021), (2) agent-based frameworks (Wang et al., 2023; Zhang et al., 2024a; b), and (3) fine-tuning-based methods (Bao et al., 2025; Zhang et al., 2024c). Among these, the fine-tuning-based solution has received the most attention due to its ability to more effectively align LLMs with the task, and our work falls into this category. In this direction, only a few studies have explored reasoning, with existing works primarily aiming to enhance explicit CoT reasoning via fine-tuning, such as RecSAVER (Tsai et al., 2024), ReasoningRec (Bismay et al., 2025), Reasoning4Rec (Fang et al., 2025), R2ec (You et al., 2025), and Exp3rt (Kim et al., ). However, these approaches rely heavily on explicit CoT, which is costly to obtain and typically leads to high inference latency. In contrast, our work introduces latent reasoning for LLM-based recommendation, which eliminates the need for CoT supervision and enables more efficient inference. As for latent reasoning, only two concurrent works (Tang et al., 2025; Zhang et al., 2025) have explored it, but neither focuses on LLM-based recommendation, and our approach to latent reasoning is fundamentally different and specifically tailored for LLM-based recommendation.
@@ 6 Conclusion
Following common practices for LLM-based recommendation works, our current experiments are limited to relatively small datasets due to the very high cost. In the future, we plan to explore the effectiveness of our method on diverse, large-scale datasets. Additionally, compared to explicit CoT, latent reasoning offers less interpretability. We plan to investigate ways to improve the explainability of latent reasoning, facilitating its wider adoption.
Ethics Statement
This work adheres to the ICLR Code of Ethics. Our research does not involve human subjects or sensitive personal data. Although we use some user interaction data, it is anonymized, publicly available, and does not contain any personally identifiable information. Moreover, we strictly follow the usage guidelines set by the dataset providers. However, when applying our method to broader scenarios in the future, it would be advisable to include additional user privacy protections, such as allowing users to control how their data is used and extending the approach to local learning settings. The experiments presented in this paper focus on algorithmic development and theoretical analysis, and do not entail any direct applications that could cause harm if misused. All authors have disclosed any potential conflicts of interest, and there are no financial or institutional relationships that may influence the objectivity of this research. The usage of LLMs is provided in Appendix K.
Reproducibility Statement
To support reproducibility, we provide a detailed description of all experimental settings, model architectures, and hyperparameters in Appendix B. Our code has been made available at https://github.com/xuwenxinedu/R3, including training scripts, evaluation scripts, and instructions for reproducing the main results. All datasets used are publicly accessible, and we provide the dataset preprocessing process in Appendix A and the preprocessing script in our code.
@@ Appendix A Dataset Pre-processing
Our experiments were conducted on data from Amazon, with primary evaluations performed on three domain-specific datasets: Toys, CDs, Games and Instruments. Following established methodology, we preprocessed each subset to achieve 5-core data integrity. To address the computational demands of LLM training while maintaining dataset validity, we implemented a dynamic temporal partitioning strategy: We iteratively process the dataset within a sliding time window, starting from October 2017 to October 2018. After applying 5-core filtering, if the resulting item count falls below 5,000, the start time is shifted back by 3 months (e.g., to July 2017). This adjustment repeats until the processed dataset contains more than 5,000 items, with the end time fixed at October 2018 throughout the process. Key dataset statistics are summarized in the accompanying table3. To align with real-world scenarios, we adopt a temporal split of the processed datasets into training, validation, and testing sets based on timestamps, following an 8:1:1 ratio. In alignment with comparative baseline requirements across all experimental conditions, we standardized the maximum sequence length at 10 for consistency in temporal pattern analysis.
@@
Instruments 66500 8312 8313 5030
Appendix B Compared Methods and Implementation Details
$\bullet$ Baselines. We consider the top-N recommendation task. To evaluate the effectiveness of our method, we compare it against both representative traditional sequential recommendation models and state-of-the-art (SOTA) LLM-based recommendation approaches:
    Caser (Tang & Wang, 2018): This is a well-known sequential recommendation approach that utilizes Convolutional Neural Networks (CNNs) to capture sequential patterns and model user preferences.
    GRU4Rec (Hidasi & Karatzoglou, 2018): This is another widely recognized method that leverages Gated Recurrent Units (GRUs) to encode sequential patterns and model user preferences.
    SASRec (Kang & McAuley, 2018): This is a highly representative sequential recommendation method that utilizes a self-attention network to model user preferences, with an architecture resembling that of the decoder-only Transformer model.
    Base: This refers to the method that directly prompts the backbone LLM to perform the recommendation task.
    COT (Tsai et al., 2024): This refers to the zero-shot version of Rec-SAVER, which uses CoT reasoning to prompt the LLM for recommendation, i.e., instructing the LLM to produce a reasoning process before generating the final results.